OpenStudy (anonymous):

Use the definition of the derivative to find the derivative of f (x) at the indicated point: f(x)=x^2-x at x=3

OpenStudy (kainui):

Show me your best guess or where you get stuck. There's no point in me explaining the entire thing when I'm sure you can do some of it already.

OpenStudy (haseeb96):

f(x)=x^2-x differentiate w.r.t x df(x)\dx= d\dx (x^2 -x) by appling power function d\dx x^n = nx^n-1 df(x)\dx = 2x -1.........eq#1 now put x =3 in eq#1 =2(3)-1 = 6-1 df(x)\dx =5

OpenStudy (haseeb96):

this is the correct answer

OpenStudy (anonymous):

I am supposed to use: f ' (x) = ( f (x+h) - f (x) )/ h as h approaches zero

OpenStudy (kainui):

Yeah, I don't know what @Haseeb96 is doing, I think he needs a nap. lol So malibu, you know what to use, show me your steps. I'll help you along, but only if you make an effort and guess.

OpenStudy (anonymous):

I think i got it, how to do it. The answer is 5.

OpenStudy (anonymous):

But I did it differently. I mean the setup too.

OpenStudy (kainui):

You should have done it with the definition, so if you mean you did it differently than Haseeb then yeah, that's great. Otherwise, it's wrong.

OpenStudy (anonymous):

Yes I used f'(x)=lim h approaches 0 f(x+h)-f(x)/h

OpenStudy (haseeb96):

then put the given x=3 in the given answer

OpenStudy (anonymous):

But thanks, @Haseeb and @Kainui.

OpenStudy (kainui):

At least you have the quick shortcut to taking the derivative so that you can check yourself! Always check yourself on "definition of a derivative" questions since it's so easy!